OpenStudy (anonymous):

a jar contains 13 red marbles, 8 blue marbles, 11 white marbles, and 2 yellow marbles. you pick one marble from the jar without looking. what is the probability of picking a red or white marble?

OpenStudy (anonymous):

I got 12/17 is this correct?

OpenStudy (nirmalnema):

for red its 13/34 and for white 11/34

OpenStudy (anonymous):

13 red marbles + 11 white marbles = 24, out of 13 + 8 + 11 + 2 = 34 total marbles. 24/34

OpenStudy (nirmalnema):

and probablity of red and white is 24/34 =12 /17

OpenStudy (nirmalnema):

@rusalka you are correct

OpenStudy (anonymous):

Thanks!

OpenStudy (nirmalnema):

welcome.